OpenStudy (anonymous):

what is the standard form of the parabola whose vertex form is y=-(x-3)^2+11 Will give medal :)

OpenStudy (anonymous):

Standard form: y = ax^2 + bx + c

OpenStudy (anonymous):

Is y=x^2+6x+2 the answer? @Blonde_Gangsta

OpenStudy (anonymous):

Yes im pretty sure

OpenStudy (anonymous):

or y=x^2+3x+11? @Blonde_Gangsta